A heater uses water from a geyser with temperature 90◦C and the atmosphere with the temperature 10◦C as reservoirs. Which is the machine's maximum efficiency?


A Carnot machine operates between reservoirs with temperatures 400 K and 300 K. At the lower temperature, the heating volume 330 is transferred J. What work does the machine do?
Answer #1

Efficiency %=( 1- T2/T1)× 100 = (1- 283/ 363)×100

= (1- 283/363)×100= 22%

T2/T1= 300/400= Q2/Q1= 330/Q1

Q1= 440J

Q1- Q2 = 440 - 330 = 100J